A 900 gram baseball is thrown horizontally at a snow pile. It enter the snow pile with a velocity of 55 m/s. It steps after travelling 6 feet inside snow pile.

a) find force of impact upon entering pile

b) what is the power exerted by the baseball upon initial impact on the snow pile if it happening in 5 milisecond
Answer #1

change in momentum = impulsive force or impact force

part a

impact force =m\Deltav=m(vf - vi)=0.900*(55-0)=49.5 N-s

ipmact =F*t =49.5

F=49.5/(5*10-3)=9.9KN

power exerted = work done by baseball=F*d=9.9KN*1.8288m    (d=6 feet=1.8288m)

power exerted=18.10512KJ